# Adaptive Random Testing with Qgrams: The Illusion Comes True

### v2

@@ 6.6. Discussion
Through theoretical analysis, simulations, and experiments on six web applications, we have demonstrated that Qgram aggregation can transform the effectiveness of ART from an “illusion” into a practical reality. While random testing or traditional ART can be effective at high failure probabilities or for less complex programs under test, the ART with Qgram aggregation techniques we have proposed have benefits for finding faults with lower failure probabilities or when programs under test are of medium to high complexity.
Qgramss vs Qgramss+i. Our results show that considering input values leads to a slightly higher efficiency, especially in low budget scenarios (i.e., AUC@20%). This might be explained by the fact that Qgramss+i tend to select slightly longer tests than Qgramss throughout the generation process (see Figure 3). On the other hand, considering method sequences only, seems to be more effective coverage-wise. Indeed, in the cases when Qgramss+i is better, the difference is small (3.6% in splittypie, and 0.85% in petclinic), while there is a larger difference when Qgramss is better (1.7% in pagekit, where the difference is statistically significant, and 10% in phoenix). Moreover, Qgramss tends to cover more unique targets than Qgramss+i on average (i.e., 1.70 vs 1.07 considering all the subjects). Although more experiments are needed to understand the impact of input values on such metrics, our results suggest that Qgramss is the preferred Qgrams technique, unless efficiency, especially at low budget regimes, is a priority.
Number of Candidates for Diversity Computation. In our experiments, we set the number of candidates to $|W|=30$, keeping it fixed throughout the test generation process. While a smaller candidate pool may be sufficient early on for covering high-probability targets–making ART with Qgram aggregation resemble Rand–a larger pool might prove beneficial later in the process for selecting test cases that reach harder-to-cover targets. Dynamically adjusting the number of candidates could improve overall coverage, a topic that future work should explore further.
@@ 6.7. Threats to validity
External validity. We evaluated our approach using a benchmark of six web applications and with specific Qgram choices, which raises questions about generalization. However, our empirical findings align with the theoretical analysis, suggesting that generalization may be feasible for scenarios that meet the theoretical assumptions behind our approach–primarily the ability to define Qgrams that allow the aggregation to be updated incrementally.
@@ 7. Conclusion and Future Work
We propose a novel framework for Adaptive Random Testing (ART) that measures diversity through an aggregation function updated incrementally. We instantiated this framework using Qgram counts as the aggregation function and entropy as the diversity measure. Our theoretical analysis shows that ART with Qgrams reduces the algorithmic cost from quadratic to linear, making the cost of diversity computation easily offset when test execution times are sufficiently long. Simulation results with a 10 $ms$ test execution time, along with empirical results from six real-world web applications, support these findings, demonstrating that ART with Qgram aggregation is more effective and efficient than both ART with distance computation and random generation, particularly for web applications with hard-to-cover targets.
In future work, we plan to explore alternative instantiations of our framework, such as using other diversity metrics like Gini impurity (Rokach and Maimon, 2005) or compression based approaches (Feldt et al., 2016). We also aim to investigate the broader applicability of our Qgram-based diversity framework beyond test generation, including its potential use in test selection, test prioritization, and for automated boundary value testing (Dobslaw et al., 2023).

### v3

@@ 6.6. Discussion
Through theoretical analysis, simulations, and experiments on six web applications, we have demonstrated that q-gram aggregation can transform the effectiveness of ART from an “illusion” into a practical reality. While random testing or traditional ART can be effective at high failure probabilities or for less complex programs under test, the ART with q-gram aggregation techniques we have proposed have benefits for finding faults with lower failure probabilities or when programs under test are of medium to high complexity.
q-gramss vs q-gramss+i. Our results show that considering input values leads to a slightly higher efficiency, especially in low budget scenarios (i.e., AUC@20%). This might be explained by the fact that q-gramss+i tend to select slightly longer tests than q-gramss throughout the generation process (see Figure 3). On the other hand, considering method sequences only, seems to be more effective coverage-wise. Indeed, in the cases when q-gramss+i is better, the difference is small (3.6% in splittypie, and 0.85% in petclinic), while there is a larger difference when q-gramss is better (1.7% in pagekit, where the difference is statistically significant, and 10% in phoenix). Moreover, q-gramss tends to cover more unique targets than q-gramss+i on average (i.e., 1.70 vs 1.07 considering all the subjects). Although more experiments are needed to understand the impact of input values on such metrics, our results suggest that q-gramss is the preferred q-grams technique, unless efficiency, especially at low budget regimes, is a priority.
Number of Candidates for Diversity Computation. In our experiments, we set the number of candidates to $|W|=30$, keeping it fixed throughout the test generation process. While a smaller candidate pool may be sufficient early on for covering high-probability targets–making ART with q-gram aggregation resemble Rand–a larger pool might prove beneficial later in the process for selecting test cases that reach harder-to-cover targets. Dynamically adjusting the number of candidates could improve overall coverage, a topic that future work should explore further.
[…]
q-grams Limitations. While q-grams automate the manual effort required for ad-hoc partitioning of the input space, they only focus on syntactic differences. A semantically meaningful partitioning would require manual, human knowledge (Zhang et al., 2016). Moreover, we acknowledge that usage of q-grams with complex data types might have limitations, which would require further investigation, particularly because q-grams are limited to local properties of the input sequence.
@@ 6.7. Threats to validity
External validity. We evaluated our approach using a benchmark of six web applications and with specific q-gram choices, which raises questions about generalization. However, our empirical findings align with the theoretical analysis, suggesting that generalization may be feasible for scenarios that meet the theoretical assumptions behind our approach–primarily the ability to define q-grams that allow the aggregation to be updated incrementally.
@@ 7. Conclusion and Future Work
We propose a novel framework for Adaptive Random Testing (ART) that measures diversity through an aggregation function updated incrementally. We instantiated this framework using q-gram counts as the aggregation function and entropy as the diversity measure. Our theoretical analysis shows that ART with q-grams reduces the algorithmic cost from quadratic to linear, making the cost of diversity computation easily offset when test execution times are sufficiently long. Simulation results with a 10 $ms$ test execution time, along with empirical results from six real-world web applications, support these findings, demonstrating that ART with q-gram aggregation is more effective and efficient than both ART with distance computation and random generation, particularly for web applications with hard-to-cover targets.
In future work, we plan to explore alternative instantiations of our framework, such as using other diversity metrics like Gini impurity (Rokach and Maimon, 2005) or compression based approaches (Feldt et al., 2016). We also aim to investigate the broader applicability of our q-gram-based diversity framework beyond test generation, including its potential use in test selection, test prioritization, and for automated boundary value testing (Dobslaw et al., 2023).
@@ 8. Data Availability
Acknowledgements.
This work was partially supported by the project Toposcope (SNF grant n. 214989).